OpenStudy (anonymous):

The surface area, S, of a sphere of radius r feet is S = S(r) = 4πr2. Find the instantaneous rate of change of the surface area with respect to the radius r at r = 4. a. 32pi b. 16pi c. 64pi d. 4pi

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):



OpenStudy (anonymous):

@iambatman @TuringTest @Abhisar @SolomonZelman @iGreen

OpenStudy (anonymous):

@iGreen please help

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

Idk

OpenStudy (michele_laino):

please, you have to calculate the first derivative oS with respecto to r, first, namely: \[\frac{ dS }{ dr }=...\]

OpenStudy (anonymous):

@Michele_Laino how would one do tht

OpenStudy (michele_laino):

what is the first derivative of r^2 with respect to r^

OpenStudy (michele_laino):

we have: \[\frac{ dS }{ dr }=\frac{ d (4 \pi r ^{2}) }{ dr }=8 \pi r\] do you agree?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so insert 4 into r and get 32pi?

OpenStudy (michele_laino):

that's right!